4. You have a stock which pays a $1.45 dividend which is expected to grow at 15% for the next five years, then at a constant 6% into the future, the discount rate is 11%. (Round Up, Two Decimals)

Answer is a. $44.77

Can you help me get to this answer?
Answer #1

D0 = $1.45

D1 = $1.45 * 1.15 = $1.6675

D2 = $1.6675*1.15 = $1.91763

D3 = $1.91763 *1.15 = $2.20527

D4 = $2.20527*1.15 = $2.53606

D5 = $2.53606 *1.15 = $2.91647

Terminal value = $2.91647 *1.06 / (0.11 - 0.06) = $61.8291

Present value = [$1.6675 / (1.11)1] + [$1.91763 / (1.11)2] +[$2.20527/ (1.11)3] + [$2.53606 / (1.11)4] + [$2.91647 / (1.11)5] + [$61.8291 / (1.11)5]

Present value = $44.77

So Stock price = $44.77